OpenStudy (anonymous):

estemate the difference inbetween them. 1. 7 7/10 2 1/10 about 5 in. about 4 in. about 6 in. about 6 1/2 in.

OpenStudy (anonymous):

is it \[7\frac{ 7 }{ 10 } and 2\frac{ 1 }{ 10 }\]

OpenStudy (anonymous):

mmmmhmmmm (:

OpenStudy (anonymous):

\[7\frac{ 7 }{ 10 } - 2\frac{ 1 }{ 10 } = 5\frac{ 6 }{ 10 }\] what 6/10 simplified? :}

OpenStudy (anonymous):

idk....lol

OpenStudy (anonymous):

don't play dumb rainbow o:< ur gonna learn something! if divide both the 6 and 10 by 2 i get what?